# Generalizations of the Squircle-Lemniscate Relation and Keplerian Dynamics

### v1

@@ 1. Introduction
and the Lamé curve $x^{4}+y^{4}=1$, popularly known as the squircle, and we provided an elementary proof of this relation. We also posed the question whether this result could be generalized to an analogous relation between Lamé curves $x^{2n}+y^{2n}=1$ for $n>2$ and some hypothetical analogs of the lemniscate. In this paper we provide generalizations of our result in [2].
@@
This equality implies that the perimeter of the sinusoidal spiral [17], given by the polar equation $r^{n}=\cos(n\theta)$ is $2^{\frac{1}{n}-1}n$ times the area of the Lamé curve [^1] G. Lamé initiated the study of these curves in 1818 [7, p. 105], hence the nomenclature. $x^{2n}+y^{2n}=1$}, [18]. We then generalize the main result of [2] relating radial sector areas of the Lamé curves to lengths of geometrically related arcs of sinusoidal spirals. We then offer a physical interpretation of this result as relating keplerian motion along Lamé curves to uniform motion along sinusoidal spirals, which we illustrate graphically in the case $n=3$. We also provide a central force interpretation of keplerian motion on Lamé curves. Finally we discuss an alternative generalization, which relates sinusoidal spirals to different generalizations of the squircle, policles whose polar equations have the form
@@ 3. The Lamé-Sinusoidal Spiral Relation
A more concise proof of Theorem 1 can be carried out using the same approach used in [5] and [11, p. 173] to prove (1). Namely evaluating the two integrals on both sides of (9) in terms of the beta $B(x,y)=\int_{0}^{1}t^{x-1}(1-t)^{y-1}\,dt$} and gamma $\Gamma(x)=\int_{0}^{\infty}t^{x-1}e^{-t}\,dt$} functions and simplifying ([1], [13]), [14], [16] yields
@@ 4. Generalized Geometric Relations
In our previous work, we utilized a result of Siegel to establish a geometric correspondence between the lemniscate and the squircle, [2, Theorem 10]. We now provide a generalized version of this theorem which relates the sinusoidal spirals to the generalized Lamé curves.
@@ 5. Physical Interpretation: Keplerian vs Uniform Motion
The relation described in Corollary 6 implies that keplerian motion along the Lamé curve $x^{2n}+y^{2n}=1$ within the polar sector $0\leq\theta\leq\frac{\pi}{4}$ corresponds geometrically to uniform motion along the sinusoidal spiral $r^{n}=\cos(n\theta)$ within the polar sector $0\leq\theta\leq\frac{\pi}{2n}$. By symmetry this correspondence extends to such a correspondence between the entire Lamé curve and the entire sinusoidal spiral, answering a question posed in [2, Remark 6].
@@ 6. General Force Law for Lamé Curves
The differential equation of an orbit under a central force $F(r)$ directed towards the origin is given by Binet’s equation [15], [3, p. 87 (3.34)]:
@@ 7. Sinusoidal Spirals and Policles
The reader might observe that Theorem 2 generalizes [2, Theorem 10], which is a reformulation of the main result [2, Theorem 1], based on a calculation of Siegel [12]. The reader might inquire whether there is a simple direct generalization of that main result relating the squircle and lemniscate to one between Lamé curves and sinusoidal spirals. Regrettably, this is not the case. However there is such a simple direct generalization if we replace the Lamé curves $x^{2n}+y^{2n}=1$ by a different generalization of the squircle: curves given by polar equations
@@
When $n=2$ this curve is a squircle, a rounded square. For general $n$ these curves look like rounded regular $2n$-gons. We call these curves policles (“polygon” ++ “circle”) by analogy with squircle (“square” ++ “circle”). With this modification, we have the following simple direct generalization of [2, Theorem 1], illustrated by the following figure for the case $n=3$.

### v2

@@ 1. Introduction
and the Lamé curve $x^{4}+y^{4}=1$, popularly known as the squircle, and we provided an elementary proof of this relation. We also posed the question whether this result could be generalized to an analogous relation between Lamé curves $x^{2n}+y^{2n}=1$ for $n>2$ and some hypothetical analogs of the lemniscate. In this paper we answer this question and provide generalizations of our result in [3].
@@
This equality implies that the perimeter of the sinusoidal spiral [18], given by the polar equation $r^{n}=\cos(n\theta)$ is $2^{\frac{1}{n}-1}n$ times the area of the Lamé curve [^1] G. Lamé initiated the study of these curves in 1818 [8, p. 105], hence the nomenclature. $x^{2n}+y^{2n}=1$}, [19]. We then generalize the main result of [3] relating radial sector areas of the Lamé curves to lengths of geometrically related arcs of sinusoidal spirals. We then offer a physical interpretation of this result as relating keplerian motion along Lamé curves to uniform motion along sinusoidal spirals, which we illustrate graphically in the case $n=3$. We also provide a central force interpretation of keplerian motion on Lamé curves. Finally we discuss an alternative generalization, which relates sinusoidal spirals to different generalizations of the squircle, policles whose polar equations have the form
@@ 3. The Lamé-Sinusoidal Spiral Relation
A more concise proof of Theorem 1 can be carried out using the same approach used in [6] and [12, p. 173] to prove (1). Namely evaluating the two integrals on both sides of (9) in terms of the beta $B(x,y)=\int_{0}^{1}t^{x-1}(1-t)^{y-1}\,dt$} and gamma $\Gamma(x)=\int_{0}^{\infty}t^{x-1}e^{-t}\,dt$} functions and simplifying ([1], [14]), [15], [17] yields
@@ 4. Generalized Geometric Relations
In our previous work, we utilized a result of Siegel to establish a geometric correspondence between the lemniscate and the squircle, [3, Theorem 10]. We now provide a generalized version of this theorem which relates the sinusoidal spirals to the generalized Lamé curves.
@@ 5. Physical Interpretation: Keplerian vs Uniform Motion
The relation described in Corollary 6 implies that keplerian motion along the Lamé curve $x^{2n}+y^{2n}=1$ within the polar sector $0\leq\theta\leq\frac{\pi}{4}$ corresponds geometrically to uniform motion along the sinusoidal spiral $r^{n}=\cos(n\theta)$ within the polar sector $0\leq\theta\leq\frac{\pi}{2n}$. By symmetry this correspondence extends to such a correspondence between the entire Lamé curve and the entire sinusoidal spiral, answering a question posed in [3, Remark 6].
@@ 6. General Force Law for Lamé Curves
The differential equation of an orbit under a central force $F(r)$ directed towards the origin is given by Binet’s equation [16], [4, p. 87 (3.34)], [2]:
@@ 7. Sinusoidal Spirals and Policles
The reader might observe that Theorem 2 generalizes [3, Theorem 10], which is a reformulation of the main result [3, Theorem 1], based on a calculation of Siegel [13]. The reader might inquire whether there is a simple direct generalization of that main result relating the squircle and lemniscate to one between Lamé curves and sinusoidal spirals. Regrettably, this is not the case. However there is such a simple direct generalization if we replace the Lamé curves $x^{2n}+y^{2n}=1$ by a different generalization of the squircle: curves given by polar equations
@@
When $n=2$ this curve is a squircle, a rounded square. For general $n$ these curves look like rounded regular $2n$-gons. We call these curves policles (“polygon” ++ “circle”) by analogy with squircle (“square” ++ “circle”). With this modification, we have the following simple direct generalization of [3, Theorem 1], illustrated by the following figure for the case $n=3$.
@@
As in Section 5, we can interpret this correspondence as relating keplerian motion around the policle to uniform motion around the sinusoidal spiral. We leave the details as an exercise for the reader.